Project

General

Profile

2015-11-17 memo-physics-report » History » Version 2

Rogers, Chris, 17 November 2015 12:56

1 1 Rogers, Chris
h1. 2015-11-17 memo-physics-report
2
3 2 Rogers, Chris
h2. 2015/03
4
5 1 Rogers, Chris
* Third draft run plan document is at http://micewww.pp.rl.ac.uk/projects/analysis/documents
6
* Lots of uncertainty around hardware, so we shall see what happens (ops meeting later)
7
* FC alignment:-
8
** Rogers has looked at settings, ~ 50 A solenoid mode looks okay (or 70 A flip mode); still not finalised
9
** Sophie is back from vacation and will start looking through "reconstruction" i.e. can we get good enough resolutions with TOF cut?
10
* Q789 switched off for TOF0 alignment
11
** nb this is a short run but should be done for TOF0 alignment
12
* SSU switched on with ECE:-
13
** Uncertainty around the hardware
14
** VB looking at trim coil currents
15
** Nb: needs some modelling of iron here - feeding in Jo Langlands (Sheffield grad student) to this effort + field mapping
16
* Straight tracks
17
** Staring at MuScat paper, looks like they were systematics dominated (errors ~ 50k muons)
18
** Working on the statistics analysis (can we get 50 k muons within momentum cut and radial cut at the absorber? what should the cuts be?)
19
20 2 Rogers, Chris
h2. 3/4 December Review
21
22 1 Rogers, Chris
* Step IV
23
** asked Ao to push through a new round of optics studies - he was distracted by PASI last week
24
** will then need to process at least one setting with full reconstruction
25
* Demonstration of ionisation cooling
26
** Baseline and baseline w/o M1 is at a reasonable point, using RMS emittance analysis
27
** CR is looking at generic phase space volume analyses (not RMS emittance); to fight the "do nothing" crowd
28
** Jaroslaw is looking at using third FC as a matching element (no news yet, Rogers pestering)
29
30 2 Rogers, Chris
h2. January optics Review
31
32 1 Rogers, Chris
* v1 agenda handed to KL, v2 attached (some familiar names)
33
* await revised ToR - but we need to start the writing up process now to have a hope to converge by January; agenda becomes the latex skeleton
34
* I am scared by the timescale
35
36 2 Rogers, Chris
h2. Step IV
37
38 1 Rogers, Chris
* Precision measurement of emittance
39
** VB working through the "TOF tracks" analysis (maybe she is in the meeting?), turning it into a reconstruction routine
40
** Awaits tracker reconstruction/geometry fixes
41
* Tracker alignment
42
** Discussion with Melissa/Jason/Durga/Tony Millington
43
** John Palin et al aligned trackers to monuments on the end face of SSU/SSD
44
** Tony Millington et al surveyed monuments on the outer rim of SSU/SSD
45
** end-face monuments are now obscured
46
** Jason Tarrant is in contact with John Palin to recover any (raw) survey data he may have, hand off to Tony
47
** Tony Millington will make a remedial survey of SSU on Wednesday this week when the diffuser goes in
48
** remedial survey on SSD may not be possible due to interference from cables and pipes
49
50 2 Rogers, Chris
h2. Cooling demo
51
52 1 Rogers, Chris
* No update on the paper, chasing on Jaroslaw "cooling demo with focus coil" optics
53
* No update on the tolerances study, will chase Alan